(Z+1) (z+5) as a Lauren series in the region 2<|z|<5


1
Expert's answer
2021-01-06T19:12:18-0500

It is not indicated around which point the expansion into Lauren series has to be done. We assume that the point is "z_0=\frac{7}{2}=3.5" . We have: "f(z)=z^2+6z+5=(z-z_0)^2+2zz_0-z_0^2+6z+5=(z-z_0)^2+(2z_0+6)(z-z_0)+2z_0^2+6z_0-z_0^2+5=(z-z_0)^2+(2z_0+6)(z-z_0)+z_0^2+6z_0+5=(z-3.5)^2+13(z-3.5)+12.25+26=(z-3.5)^2+13(z-3.5)+38.25"

The answer is: "f(z)=(z-3.5)^2+13(z-3.5)+38.25"